OpenStudy (anonymous):

PLEASE HELP!!!! GIVING OUT MEDALS. List the possible rational roots of P(x) given by the Rational Root Theorem. P(x)= 3x^4-4x^3-x^2-7

OpenStudy (anonymous):

take all fractions where the numerator divides \(7\) and the denominator divides \(3\) don't forget the \(\pm\)

OpenStudy (anonymous):

sorry can you go a little more into depth?

OpenStudy (anonymous):

the divisors of \(7\) are only \(1\) and \(7\) the divisors of \(3\) are \(1\) and \(3\) \[\pm1,\pm 7,\pm\frac{1}{3},\pm\frac{7}{3}\]

OpenStudy (anonymous):

ohh okay. why do you choose 7 and 3?

OpenStudy (anonymous):

\(7\) is the constant, the leading coefficient is \(3\)

OpenStudy (anonymous):

oh awesome thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@satellite73 is there anyway you can look at this one for me too please? Find any rational roots of P(x). P(x)=x^3+5x^2+x+5